00:52 Sending OSCO preset
  • Seeing 0.9" on GCS
  • Mastrerdiagnostic process went down during pupil check. Cycled WFS Control gui, rebinned and verified all happy and communicating, and resending preset
  • 1:00 OSCO preset reset.
  • Failed to apply optical gail. Failed to stabilize in 10 iterations. Looks like sky is variable with cloud cover and twilight.
  • Josh resent the preset in intervention mode. He had cancelled the preset in the flao gui, but the loop did not open when preset resent so had to manually open from WFS gui.
  • 1:10 OSCO resent in intervention with cloud and high winds, 1.7m/s on the secondary art your northerly pointing
    Gains 1.01, 0.19, 0.17 with optical gain of 0.415
  • No NCPA luci images luci1.20201106.0003-0022
    tn#20201106_011546, tn#20201106_011753
  • With NCPA on luci1.20201106.0023-41, optical gain 0.37
    tn#20201106_012226, tn#20201106_012456
00:29 Slewing to Massi field
  • Seeing 1.1" on the DIMM
  • OCAM mag 14.9
    Bin 3, Gains 879, OCAM 600
  • Seeing has blown up to 2.1". Set 10 modes and are waiting this out.
  • 1:43 Reoptimizing, 0.128 optical gain, tt0.63, ho1=0.10, h02=0.34, NCPA on, DIMM 1.0
  • LUCI did not recognize that we completed AO preset. We aborted the LUCI preset, skipped first item. Sent zero, zero abs offset through LUCItelsvc to verify luci and AO talking and resuming science script at second line item
  • seeing highly variable 1.2 -1.9
    tn#20201106_015229
    paused science, but GOpt is working very well, no skipped frames.
  • Winds up to 2.5m/s on the adsec so still ok
  • Clouds followed by and explosion in the seeing. 2.6+". On target with 10 modes for now
  • 2:17 Dimm 1-1.17". Reoptimized. After second reoptimization there is no way to reapply optical gain through FLAO gui. So done through GOPT gui only.
    1. Renoramlize Loop Gains
    2. Apply to Gopt
    3. Apply to Slope Computer
  • 2:26. Seeing has clown up again to 1.62" on the DIMM
2:26 Slewing to Standard. Clouds and highly variable seeing. SL preset.

2:30 Closing for high winds. Winds have remained below 2.9m/s on the adsec.

10:02. Josh is calling the night. Warming OCAM and shutting down.

-- JenniferPower - 06 Nov 2020
Topic revision: r2 - 06 Nov 2020, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback